OpenStudy (anonymous):

Solve the triangle. A = 52°, b = 10, c = 7 a ≈ 12, C ≈ 47.9, B ≈ 80.1 a ≈ 7.9, C ≈ 43.9, B ≈ 84.1 No triangles possible a ≈ 12, C ≈ 43.9, B ≈ 84.1

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Use the cosine rule: \(a^2=b^2+c^2-2bc~ cos(A)\) after that use sine rule to solve for the remaining angles: \(\dfrac{sin(A)}{a}=\dfrac{sin(B)}{b}=\dfrac{sin(C)}{c}\)

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

Mathmate is correct, all you have to do is take the equation and replace the values.

OpenStudy (anonymous):

but i have down syndrome can u tell me the answer

OpenStudy (owlcoffee):

I will not give you an answer, but I will help you out. Let's take the cosine rule and replace the values: \[a^2 = (10)^2 + (7)^2 -2(10)(7)Cos(52^{o})\]

OpenStudy (anonymous):

100+49-70cos(52)

OpenStudy (anonymous):

?

OpenStudy (owlcoffee):

yes, now operate all that.

OpenStudy (anonymous):

-52.19

OpenStudy (anonymous):

?

OpenStudy (owlcoffee):

no, it should be: \[a^2 = 105.903\]

OpenStudy (owlcoffee):

solve for "a"

OpenStudy (owlcoffee):

square root of 105.903

OpenStudy (anonymous):

10.2

OpenStudy (owlcoffee):

good, that's the side "a", now to use the sine rule to find the remaining angles.

OpenStudy (anonymous):

thats not even an answer choice

OpenStudy (anonymous):

@TheSmartOne